A ball is thrown upward from a height of 15 feet with an initial velocity of 5 ft/s. use the formula h(t) = -16t^2 + 5t + 15 what is the maximum height the maximum height of the ball?
Answer by lwsshak3(11628)

A ball is thrown upward from a height of 15 feet with an initial velocity of 5 ft/s. use the formula h(t) = -16t^2 + 5t + 15 what is the maximum height the maximum height of the ball?
***
h(t) = -16t^2 + 5t + 15
complete the square:
h(t) = -16(t^2 - (5/16)t+(25/1024))+25/64 + 960/64
h(t) = -16(t- (5/32))^2)+985/64
This is an equation of a parabola that opens downward with vertex at (5/32,985/64)
maximum height of the ball=985/64≈15.39 ft